A farmer decides to enclose a rectangular garden, using the side of the barn as one rectangle. What is the maximum area that the farmer can enclose with 80ft. Of fence. What should the dimensions of the garden be to give this area?
Answer by lwsshak3(11628)

A farmer decides to enclose a rectangular garden, using the barn as one side of the rectangle. What is the maximum area that the farmer can enclose with 80ft. Of fence. What should the dimensions of the garden be to give this area?
let x=length of rectangle (length of barn)
(80-x)/2=width of rectangle
Area=length*width=x(80-x)/2=80x-x^2=40x-x^2/2
A=-x^2/2+40x
complete the square:
A=-1/2(x^2-80x+1600)+800
-1/2(x-40)^2+800
This is an equation of a parabola that opens down with vertex at (40,800)
dimensions of the garden: 40 by 20 ft